November 1996  ICTCM IX (Reno)  5

Example: The Parachute Problem

Edwards and Penney (Prentice Hall, 1996, p. 95)

A woman bails out of an airplane at an altitude of 10,000 ft, fall freely for 20 s, then opens her parachute. How long will it take her to reach the ground? Assume that ρ = 0.15 without the parachute and that ρ = 1.5 with the parachute. (Suggestion: First determine the height and velocity when the parachute opens.)

NOTE: ρ = \frac{k}{m}